### v1

# Repeating Flares, X-ray Outbursts and Delayed Infrared Emission: A Comprehensive Compilation of Optical Tidal Disruption Events

[caption]
Repeating Flares, X-ray Outbursts and Delayed Infrared Emission:
A Comprehensive Compilation of Optical Tidal Disruption Events
Tidal disruption events (TDEs) have been proposed as valuable laboratories for studying dormant black holes. However, progress in this field has been hampered by the limited number of observed events. In this work, we present TDECat, a comprehensive catalogue of 134 confirmed TDEs (131 optical TDEs and 3 jetted TDEs) discovered up to the end of 2024, accompanied by multi-wavelength photometry (X-ray, ultraviolet (UV), optical, and infrared) and publicly available spectra. We also study the statistical properties, spectral classifications, and multi-band variability of these events. Using a Bayesian Blocks algorithm, we determine the duration, rise time (), decay time (), and their ratio for 103 flares in our sample. We find that these timescales follow a log-normal distribution. Furthermore, our spectral analysis shows that most optical TDEs belong to the TDE-H+He class, followed by the TDE-H, TDE-He, and TDE-featureless classes, which is consistent with expectations from main sequence star disruption. Using archival observations, we identify four new potentially repeating TDEs, namely AT 2024pvu, AT 2022exr, AT2021uvz, and AT 2019teq, increasing the number of known repeating events. In both newly identified and previously known cases, the secondary flares exhibit a similar shape to the primary. We also examine the infrared and X-ray emission from the TDEs in our catalogue, and find that 14 out of the 18 infrared events have associated X-ray emission, strongly suggesting a potential correlation. Finally, we find that for three subsamples (repeating flares, infrared-emitting events, and X-ray-emitting events), the spectral classes are unlikely to be randomly distributed, suggesting a connection between spectral characteristics and multi-wavelength emission. TDEcat enables large-scale population studies across wavelengths and spectral classes, providing essential tools for navigating the data-rich era of upcoming surveys such as the Legacy Survey of Space and Time.
Key Words.:
1 Introduction
A tidal disruption event (TDE) occurs when a star ventures too close to a supermassive black hole (SMBH), where differential gravitational forces exceed the star’s self gravity, causing it to be torn apart by tidal forces. This occurs only if the tidal radius, , which is the critical radius within which an object is disrupted by the SMBH, is larger than the event horizon. There is a theoretical upper limit to the SMBH’s mass known as the Hills mass, since . For a solar mass star, this limit is approximately (Hills 1975). About half of the stellar material remains bound to the SMBH forming an accretion disc, while the other half is ejected (Rees 1988). This process produces a characteristic flare with a sharp rise in brightness, followed by a gradual decay leading to a plateau. These flares can be observed across the electromagnetic spectrum, from radio to hard X-rays and typically exhibit broad hydrogen and helium lines. The expected detection rate of optical TDEs has been estimated at a few galaxy-1 yr-1 (Yao et al. 2023). Over the past half decade, approximately 10-20 new optical TDEs have been detected annually.
The first TDE was identified during the ROSAT all-sky X-ray survey, with the observed soft X-ray emission attributed to a newly formed accretion disc (Bade et al. 1996; Grupe et al. 1999; Saxton et al. 2020). Since then, numerous TDEs have been detected by various instruments and surveys. In the X-ray band, both XMM-Newton and eROSITA have observed multiple TDEs and candidates. The Chandra X-ray Observatory and the BAT telescope aboard the Neil Gehrels Swift Observatory (Roming et al. 2005; Burrows et al. 2005) have played important roles in follow-up X-ray observations of TDEs discovered at optical wavelengths. Additionally, Swift’s UVOT telescope has enabled observations of these events in the ultraviolet.
Optical surveys have significantly contributed to our understanding of TDEs. The All-Sky Automated Survey for Supernovae (ASAS-SN; Shappee et al. 2014; Kochanek et al. 2017; Hart et al. 2023), a global network of small automated telescopes scanning the entire sky for supernovae and other transients, has identified numerous TDEs through their distinctive optical flares. Similarly, the Zwicky Transient Facility (ZTF; Graham et al. 2019; Bellm et al. 2019; Masci et al. 2019), with its high-cadence, wide-field survey capabilities, has played a key role in the discovery of optical TDEs, by rapidly surveying vast areas of the sky. The Asteroid Terrestrial-impact Last Alert System (ATLAS; Tonry et al. 2018; Heinze et al. 2018; Smith et al. 2020; Shingles et al. 2021) has also been crucial to long term monitoring of TDEs by providing photometric data in two optical wavebands.
The Gaia photometric Science Alerts system (GSA111http://gsaweb.ast.cam.ac.uk/alerts Hodgkin et al. 2021) has also contributed to TDE discoveries by detecting and providing optical photometry of transient events. Several other surveys were instrumental in early TDE observations, including the Catalina Real-Time Transient Survey (CRTS; Drake et al. 2009), Lincoln Near-Earth Asteroid Research (LINEAR; Stokes et al. 2000), Panoramic Survey Telescope and Rapid Response System (Pan-STARRS; Chambers et al. 2016; Waters et al. 2020; Magnier et al. 2020a, c, b; Flewelling et al. 2020), the Palomar Transient Factory (PTF Law et al. 2009; Rau et al. 2009) and its successor, the Intermediate Palomar Transient Factory (iPTF222https://www.ptf.caltech.edu/iptf/).
Beyond optical wavelengths, TDEs have been observed in the infrared, initially by the Wide-field Infrared Survey Explorer (WISE; Wright et al. 2010) and later by the Near-Earth Object WISE Reactivation (NEOWISE; Mainzer et al. 2011, 2014). Spectroscopic follow-up observations, including the Sloan Digital Sky Survey (SDSS; York et al. 2000) and the Palomar 200-inch Hale Telescope, have been crucial in classifying TDEs based on their distinctive spectral signatures.
Despite the growing numbers of TDEs no comprehensive catalogue currently exists. Such a resource would enable robust statistical analyses by providing a dataset large enough to conduct meaningful studies, even if the sample is incomplete. The primary goal of this paper is to compile and present all publicly available photometric and spectroscopic data for confirmed optical TDEs up to the end of 2024. Additionally, we investigate different sub-categories of optical TDEs, including those displaying repeating flares, X-ray outbursts, and infrared emission.
Our paper is organized as follows: In Sect. 2, we outline the construction of our main sample of confirmed optical TDEs, along with a sample of TDE candidates. Section 3 describes the data collection process for the catalogue, while Sect. 4 delves into the properties of our sample. Finally, in Sect. 5 we discuss our results and their implications, followed by a summary in Sect. 6.
2 Sample selection
Our sample consists of events identified either in the Transient Name Server (TNS333https://www.wis-tns.org/; 105 sources) or in the literature (29 sources). We differentiate between confirmed TDEs and TDE candidates by organizing them into two separate catalogues. The main catalogue consists of TDEs that can be found at least in one photometric survey and have an available classification spectrum from the time of the flare. We also include TDEs with featureless spectra (see Sect. 4.2) and the three widely accepted jetted TDEs from the literature. TDEs that do not meet these criteria are designated as TDE candidates and are included in a separate TDE candidates catalogue.
2.1 Main catalogue
The first step in constructing our sample was to retrieve all objects classified as TDEs from TNS. At the time of this study, TNS listed 98 classified TDEs up to the end of 2024. However, AT 2018meh is the same event as AT 2023clx, so we include only AT 2023clx, reducing the TNS-TDE count to 97. From TNS, we also include 4 additional TDE-H+He events, 3 TDE-He events (see Sect. 4.2 for spectral classification) and a TDE-H+He event that is classified in an AstroNote (AT 2024ule444https://www.wis-tns.org/astronotes/astronote/2024-318). Hence, our TNS-TDE sample consists of a total of 105 transients.
Beyond TNS, additional TDEs have been identified in various published sample studies. Hammerstein et al. (2023) present a sample of 30 TDEs, observed during the first phase of the ZTF survey (see their Table 1). Of these, 12 are not classified in TNS: AT 2018lni, AT 2018jbv, AT 2019cho, AT 2019mha, AT 2019meg, AT 2020ddv, AT 2020ocn, AT 2020opy, AT 2020mbq, AT 2020qhs, AT 2020riz and AT 2020ysg. We note that this sample includes almost all the TDEs from the sample of van Velzen et al. (2021), except for AT 2019eve. This transient showed spectral and light curve evolution that made its initial TDE classification ambiguous (Hammerstein et al. 2023). For this reason, we opt to include AT 2019eve in Sect. 2.2, where we present strong TDE candidates.
Another TDE sample is presented in Yao et al. (2023) which lists 33 TDEs (see their Table 3). Of these, only 3 are absent from both the Hammerstein et al. (2023) sample and the TNS-TDE sample, namely: AT 2019baf, AT 2019cmw and AT 2020abri. While these 3 were classified as TDEs in Yao et al. (2023), there are uncertainties regarding their classification. AT 2019baf also appears in Somalwar et al. (2023a) as a TDE but it lacks a classification spectrum from the time of the flare. Additionally, Somalwar et al. (2023a) cite an unpublished work as the basis for its classification. A similar situation applies to AT 2019cmw, which is also referenced in an unpublished paper in Yao et al. (2023). AT 2020abri, meanwhile, has an optical spectrum taken 395 days after the peak of its optical flare, well after the event had likely faded. Its classification is based on: 1) persistent blue colour and lack of cooling, which is inconsistent with the majority of SNe and 2) the combination of weak H emission and strong H absorption, which indicate that the host galaxy is post-starburst (where TDE rates are enhanced; see Sect. 4.3). Since none of these three sources have spectra from the time of their flares, we exclude them from the main sample and place them in the TDE candidates sample.
Additionally, several studies focus on individual TDEs that are neither classified in TNS nor included in large sample studies (e.g., van Velzen et al. 2021; Hammerstein et al. 2023; Yao et al. 2023). We include 17 such sources in our catalogue, with detailed descriptions provided in Appendix A.
In total our main catalogue consists of 134 TDEs (131 optical TDEs and 3 jetted TDEs), including all confirmed events up to 2024. As it was briefly mentioned in the Introduction, the creation of a catalogue of all the known optical TDEs so far can allow statistical works, which were previously unable to be carried out due to small sample sizes. We note that this is not a complete sample, since many detected events remain unclassified and certain subtypes (e.g., events detected in infrared) are not fully explored.
2.2 TDE candidates
Alongside our main TDE sample, numerous transients have been classified as TDE candidates. These sources are included in a supplementary table, available on our GitHub page (see Sect. 3). The setup of this file is shown in Table 1. The first four columns list the name and coordinates, while the fifth and sixth columns include remarks on the candidate status and reference studies, respectively. To compile the TDE candidates sample, we utilized TDExplorer555https://jminding.pythonanywhere.com/main/tdes-by-name, which is a catalogue of TDEs and candidates identified through natural language processing applied to the abstracts of papers.
AT Name | Alternative Names | RA | DEC | Comments | Reference | ||
- | OGLE16aaa | 01:07:20.88 | -64:16:20.70 | TDE candidate in a weak AGN | Wyrzykowski et al. (2017) | ||
AT 2019eve | ZTF19aatylnl/Gaia19bti | 11:28:49.650 | +15:40:22.30 | Persistent line one year post peak. |
AT 2021lwx | ZTF20abrbeie | 21:13:48.405 | +27:25:50.46 | Ultraluminous, long duration transient | Subrayan et al. (2023) | ||
AT 2024kmq | ZTF24aapvieu | 12:02:37.273 | +35:23:35.22 | Luminous, fast, red transient | Ho et al. (2025) | ||
- | Swift J1112+82 | 11:11:47.32 | -82:38:44.20 | Likely jetted TDE | Brown et al. (2015) |
3 Data
The catalogue is available online on a dedicated GitHub666https://github.com/dlangis/TDECat page. On this page we have compiled all publicly available photometric and spectroscopic data for the TDEs in our sample. The full catalogue is also available via a local Python-based app. Below we summarize how and from where we collected the photometric and spectroscopic data included in the catalogue.
3.1 Optical and Infrared Photometry
For compiling the optical and infrared photometric data, we used the Black Hole Target Observation Manager (BHTOM777https://bhtom.space), a web server designed to provide astronomers easy access to astronomical data and a network of telescopes. One of the key features of BHTOM allows the user to compile all the available, archived photometric data in an interactive plot and a .csv downloadable file. The catalogue includes data from several surveys, namely LINEAR, CRTS, ZTF, iPTF, SDSS, ASAS-SN, ATLAS, NEOWISE, Pan-STARRS and GSA.
We also manually searched for CRTS light curves of LSQ12dyw from the Catalina Surveys Data Release 2888http://nesssi.cacr.caltech.edu/DataRelease/ (Drake et al. 2009). Furthermore, for specific TDEs, we obtained archival photometric data from previously published studies:
Since different surveys use different photometric methods, we converted all magnitudes and estimated flux densities in the AB system (Oke & Gunn 1983) throughout this work. The monochromatic AB magnitude is defined as , where is the spectral flux density and is the zero-point. ZTF magnitudes are calibrated using a source with colour g - r = 0 in the AB photometric system. ATLAS, iPTF, Pan-STARRS and ASAS-SN band photometry is provided in the AB system. On the contrary, CRTS, ASAS-SN band, Swift and NEOWISE999Infrared zero-points for conversion; https://wise2.ipac.caltech.edu/docs/release/allsky/expsup/sec4_4h.html photometry is given in the Vega system. For Vega-to-AB magnitude conversions, we used the values presented in Table 1 of Blanton & Roweis (2007). The Gaia survey uses an internal calibration different from both the AB and Vega systems.
3.2 X-ray Photometry
To study the X-ray emission from TDE sources we used data from Swift, Chandra and XMM-Newton archives. For each TDE source, we analysed all observations available from these missions, and evaluated X-ray fluxes and spectra for each observation. In addition to Swift-XRT, Chanra-ACIS and XMM-Newton-EPIC data, we searched for X-ray counterparts in the 13th data release of the fourth XMM-Newton serendipitous source catalogue (4XMM-DR13, Webb et al. 2020) and eROSITA main catalogue (eRASS1, Merloni et al. 2024). A detailed description of the X-ray data reduction is provided in Appendix B.
3.3 Ultraviolet Photometry
To evaluate TDE variability in UV-optical bands, we used data from the Swift Ultraviolet and Optical Telescope (UVOT) (Roming et al. 2005), which provides photometry in three near-ultraviolet (UVW2, UVM2 and UVW1) and three optical (U, B, V) bands.
UVOT data were downloaded from HEASARC101010https://heasarc.gsfc.nasa.gov/ and reduced using standard procedures111111http://www.swift.ac.uk/analysis/uvot/image.php (HEAsoft package v. 6.33.2). After checking the correct World Coordinate System alignment with USNO-B Catalogue (Monet et al. 2003), we combined image extensions using uvotimsum and merged exposure maps.
Sources were detected using uvotdetect. As with XRT data (see Appendix B.1), we assigned UVOT counterparts based on proximity to the TDE source coordinates: if a UVOT sources was detected within , its coordinates were adopted. If no source was detected within , we used the TDE sources coordinates.
We performed source photometry using uvotsource with an aperture radius of for all filters. The background region was an annulus centred at the source with inner and outer radii of and , respectively, removing emission from overlapping sources. All magnitudes were corrected for Galactic extinction using reddening estimates from Schlafly & Finkbeiner (2011) and the extinction model from Fitzpatrick & Massa (2007). We note that these data are not host subtracted.
3.4 Spectroscopic data
All optical spectra included in the catalogue were obtained from either TNS or the Weizmann Interactive Supernova Data Repository121212https://www.wiserep.org (WISeREP; Yaron & Gal-Yam 2012). Hence, the catalogue does not include optical spectra for TDEs not classified in TNS. For the 12 TDEs from Hammerstein et al. (2023), the classification spectra can be found in their Figs. 2 and 14-16. For AT 2018lni, AT 2019cho, AT 2019mha and AT 2019meg, the classification spectra can be found in van Velzen et al. (2021).
Additionally, we list references for (mainly optical) spectra of 13 TDEs that are not in the TNS-TDE sample or included in a previously published catalogue.
AT 2023vto: Fig. 4 in Kumar et al. (2024)
AT 2022agi: Figs. 2 (optical), 3 (UV) in Sun et al. (2024)
AT 2017eqx: Figs. 1, 3 in Nicholl et al. (2019)
LSQ12dyw†131313† signifies the sources that have spectra available in WISeREP.
PTF09djl†: Figs. 4, 14 in Arcavi et al. (2014)
PTF09ge†: Fig. 12, 14 in Arcavi et al. (2014)
PS1-10jh†: Fig. 1 in Gezari et al. (2012)
iPTF15af†: Figs. 3 (optical), 4 (UV) in Blagorodnova et al. (2019)
iPTF16axa†: Figs. 7, 8 in Hung et al. (2017)
ASASSN-14li†: Fig. 3 in Holoien et al. (2016b)
ASASSN-14ae†: Figs. 4, 5 in Holoien et al. (2014)
ASASSN-15oi†: Fig. 3 in Holoien et al. (2016a)
4 Exploring the catalogue
After having constructed our main sample, we now explore its various subcategories and interesting objects. Specifically, we present the statistics and spectral classes of the TDEs in the catalogue, along with new results on TDEs with repeating flares, delayed infrared emission and X-ray outbursts.
4.1 Catalogue statistics
In this section we analyse the flaring characteristics of the TDE population, in terms of rise/decay times and the event durations. To achieve this, we apply the Bayesian Blocks (Scargle 1998; Scargle et al. 2013) algorithm which partitions a one-dimensional flux time-series into blocks, each modelled by a constant flux. To avoid over-fitting, we include a user-specified penalty parameter for each additional block.
Given flux measurements and their uncertainties , we first compute cumulative weighted sums (with ) up to each index: . These cumulative sums allow efficient computation of sums over any sub-interval via . Consequently, the maximum-likelihood estimate for the constant flux in a block spanning is .
Assuming Gaussian errors, the log-likelihood of modelling data in the interval with a single constant value is
By expanding the summation in the exponent, one obtains
Moreover, the algorithm uses dynamic programming to determine the optimal segmentation. For each index (ranging from 1 to the total number of data points ), all possible previous boundaries (from to ) are considered. We compute:
where
is the optimal (maximum) log-likelihood score obtained for the segmentation of the interval ,
is the log-likelihood for modelling the new block with constant flux, and
penalty is a user-specified constant subtracted each time a new block is added. We chose a default value of 10 for the penalty to avoid over-fitting or under-fitting (i.e. small or large penalty values respectively).
The optimal segmentation is obtained by choosing the that maximizes candidate_score followed by a backtracking step (from to ) to recover the optimal boundaries.
Applying the Bayesian Blocks algorithm to the flux light curves yields blocks within which the flux is assumed to be constant. For each block, we record its mean flux value. The peak of the flare is then identified as the block with the highest mean flux. We primarily use ZTF(zg) light curves (76/103). In case ZTF data are not available or the ZTF(zg) light curve is poorly sampled, we use the light curve with the best sampling. This results in 11 ZTF(zr) light curves, 6 ATLAS light curves, 5 ASAS-SN light curves, 1 GSA(G) light curve, 1 PS1 light curve, 1 CRTS(CL) light curve and 2 PTF light curves.
To estimate the rise and decay times of the flare, we apply a threshold relative to the peak flux. Starting from the peak block, we move to lower flux levels (to lower or higher block indices) until encountering a block with a mean flux below this threshold. The first sub-threshold block on the left marks the start of the rise, while the first sub-threshold segment on the right marks the end of the decay. From these boundaries, we calculate the rise time, , and decay time, , and determine their ratio .
An example of this process is displayed in Fig. 1. For this example, the rise time is 25698 days, the decay time 582192 days and the peak time 5980322 MJD. The duration of the flare, calculated as the sum of the rise and decay times, is roughly 838 days.
We applied this method to 103 TDE flares in our sample, excluding most of the sources detected in 2024 and some older ones where rise or decay times could not be reliably estimated due to large gaps in the light curves. We also included both flares for AT 2022dbl and AT 2020vdq (see Sect. 4.3). Figure 2 shows the distribution of the common logarithm (i.e. logarithm with base 10) of the ratio in our sample. A normal distribution fit to this data yields and mean of -0.49, with a Kolmogorov-Smirnov (K-S) test p-value of 0.94 (K-S statistic = 0.05). This means that the ratio distribution is consistent with a log-normal distribution.
Figure 3 displays the duration, rise and decay times of the flares. Gaussian fits to these distributions yield means of 35020 days, 845 days and 25020 days for the duration, and (the error is the standard error of the mean) and standard deviations of 207 days, 55 day and 167 days respectively. The p-values are 0.91, 0.85 and 0.70 (K-S statistic of 0.054, 0.059 and 0.068) for the durations, and respectively, consistent with log-normal distributions.
The bottom right panel of Fig. 3 shows a scatter plot of with error bars calculated as , where is half the width of the rise/decay blocks and is half the width of the peak block. Given errors in both the independent () and dependent () variables, we employ the orthogonal distance regression (ODR) to fit a best-fit line while accounting for both uncertainties. The best-fit ODR line is shown as the red solid line in the bottom right panel of Fig. 3, while the grey contour indicates the 95% confidence interval. It is evident that there is large scatter around the best-fit line, , where and , with a 95% confidence interval of 0.8374 to 1.1227 and -0.6859 to -0.0168 for and respectively. We tested whether the correlation between and is real using a t-test, the Spearman correlation coefficient (, with a corresponding p-value ), and bootstrapping. In all our tests we cannot reject the null hypothesis of a strong, positive correlation between the rise and the decay times.
Figure 4 shows the distribution of the redshift for the TDEs in the main sample. Interestingly, this distribution also follows a normal distribution in log-space with equal to the mean and the standard deviation of the data respectively (K-S statistic equal to 0.0579 and ). We discuss this further in Sect. 5.1.
Additionally, we use the Anderson-Darling (AD) normality test to further examine the validity of the log-normal behaviour of the timescale and redshift distributions. In all cases, we cannot reject the null hypothesis that the distributions are normal (in log-space).
4.2 Spectral classification
In this section we investigate the different spectral classes of the TDEs in our sample. For consistency, we adopt the three spectral classes defined in van Velzen et al. (2021), along with the TDE-featureless class introduced by Hammerstein et al. (2023):
TDE-H, where the spectrum exhibits distinct and broad H and H emission lines
TDE-H+He, where the spectrum shows both broad H and H emission lines and broad He ii emission features
TDE-He, where the only distinct broad feature in the spectrum appears near the He ii emission line with no detectable Balmer emission lines
TDE-featureless, where the spectrum primarily displays host absorption lines with no distinct emission features characteristic of the other three classes
We note that for most objects in our sample only the classification spectrum is available, typically from the time of the flare. This makes the spectral classification uncertain, as the spectral properties of some TDEs evolve over time (e.g. Charalampopoulos et al. 2022). Several previous studies have already classified the TDEs in their samples, including van Velzen et al. (2021), Hammerstein et al. (2023), and Yao et al. (2023). In some cases, the resulting classifications are mixed. For example, AT 2019mha, AT 2019bhf, and AT 2018hyz appear in two or even all of the above studies, but have different classifications. Additionally, Charalampopoulos et al. (2022) provided spectral classifications for the TDEs in their sample (see their Table 4), some of which contradict the classifications assigned by van Velzen et al. (2021) for overlapping sources (namely ASASSN-15oi and PTF09ge). Furthermore, Charalampopoulos et al. (2022) identified three TDEs: AT 2018hyz, AT 2017eqx and ASASSN-14ae, that show spectral evolution over time (see their Table 4). Throughout this analysis, we assign events to the TDE-H+He class if they have been classified as such in at least one observation epoch. Additionally, 6 more transients listed in TNS have already been assigned a spectral class (3 TDE-H+He and 3 TDE-He). Moreover, spectral classifications or information on the broadness of the emission lines are available in TNS AstroNotes or classification reports of several TDEs. Finally, since the classification spectra are publicly available for almost all sources in the catalogue, we classify the remaining TDEs in this work.
The results of this analysis are presented in Fig. 5. To summarize, 17.6% (23) of the TDEs belong to the TDE-H class, 12.2% (16) are classified as TDE-He, 60.3% (79) fall into the TDE-H+He class and 9.9% (13) are TDE-featureless. We note that even when considering TDEs with multiple classifications based on spectra from different epochs, the overall distribution of spectral classes remains largely unchanged. If we consider the alternative classifications (PTF09geTDE-He, ASASSN-14aeTDE-H, ASASSN-15oiTDE-He, AT 2017eqxTDE-He, AT 2018hyzTDE-H, AT2019bhfTDE-H, AT2019mhaTDE-H), the aforementioned percentages change to 20.6% (27), 13.7% (18), 55.7% (73) and 9.9% (13) for the TDE-H, TDE-He, TDE-H+He and TDE-featureless classes. The spectral classification of each TDE, along with the corresponding references, can be found on the catalogue GitHub page. We discuss the implications of our results in Sect. 5.2
4.3 Repeating flares
A few TDEs in our catalogue exhibit repeating flares. These flares may correspond to separate disruption events, where different stars are disrupted each time, or they may result from the partial disruption of the same star. In some cases, the re-brightening could be caused by emission from an accretion disc, toward the later stages of the flare. Additionally, AT 2021mhg represents a unique case believed to be a TDE followed by a supernova (SN; Somalwar et al. 2023b). Below, we present our sample of repeating-flare TDEs and our analysis of four previously unreported cases: AT 2024pvu, AT 2022exr, AT 2021uvz, and AT 2019teq. To the best of our knowledge, these TDEs have not yet been reported as exhibiting a repeating flare. We also study AT 2020acka and AT 2019ehz, who have been recently referenced briefly for having re-brightening bumps (Guo et al. 2025; Zhong 2025).
In total we identify 12 TDEs with multiple flares. The corresponding light curves highlighting the repeating outbursts are shown in Fig. 6. AT 2018fyk is plotted separately in Fig. 11, since it is the only TDE with Gaia photometry, which cannot be converted to flux. We select the best sampled light curves from the available surveys to maximize the visibility of the flares, plotting only the relevant flare regions.
We approximate the host galaxy emission as the average flux measured prior to the flare. To test the validity of this assumption, we split the light curves into bins corresponding to different ZTF observational seasons for each TDE, compute their average flux and standard deviation and compare the averaged values to the pre-flare flux using a test.
In all cases, the reduced values and corresponding p-values are consistent with the flux being constant, indicating that the pre-flare light curves can be approximated by the average flux. This average flux is then subtracted from the entire light curve to remove the host galaxy’s contribution, retaining only the positive flux variations associated with the transient event. The host-subtracted light curves are displayed in Fig. 6.
The bottom panels of Fig. 6 illustrate the multiple flares of each TDE, overlaid for direct comparison. We normalize the flux to the corresponding flare peak for comparison. We observe that the TDEs in the repeating flare subsample display flares that are similar in shape with the confirmed TDE flares.
In the following, we examine in detail the repeating flare TDEs that have not previously been studied extensively, namely: AT 2024pvu, AT 2022exr, AT 2021uvz, AT 2020acka, AT 2019teq and AT 2019ehz. The remaining repeating flare TDEs are discussed in Appendix C.
4.3.1 AT 2024pvu
AT 2024pvu showed a previous outburst approximately 17.9 years prior to the 2024 flare. The first flare was detected by the CRTS, while the second, most recent flare was observed by the ZTF survey in all three available bands (, and ). The two flares are plotted in the upper panel of Fig. 6(a), while their shapes are compared in the bottom panel. A visual inspection, reveals that the two flares exhibit similar shapes.
Unfortunately, due to limited photometric and spectroscopic data of the initial flare, we cannot confirm whether it was caused by a TDE or another type of transient. Furthermore, there are no spectra available for the host galaxy, or any additional information indicating that it is a post-starburst galaxy, where as we discuss later, TDE rates are significantly enhanced. However, we can confidently exclude the possibility that both flares resulted from the disruption of the same star. If we assume a solar-like star, a SMBH with a mass of and an orbital period of 17.9 years, an orbital eccentricity of 0.99932 would be required. Such an orbit would be highly unstable, making this scenario unlikely.
4.3.2 AT 2022exr
AT 2022exr displays one of the most unique light curves in our sample, with a distinct double peak, followed by a third, smaller peak as (see in the upper panel of Fig. 6(b)). The ZTF -band light curve highlights in detail all three flares, even though there is a gap in coverage during the rise of the first flare. All three flares occurred within 250 days, with the second flare appearing before the first one had fully faded and the third flare before the second had ended.
Since no publicly available spectra exists for the second and third flares, we cannot confirm their TDE nature. However, likely all three flares are related to TDE activity. All three normalized light curves exhibit a similar shape, albeit with different peak fluxes. As we discuss in more detail in Sect. 5, this feature is characteristic of repeating TDE related flares.
We assume that all three flares are TDE-related, as it is highly unlikely that they are caused by separate transient events. We can study the partially-disrupted TDE (pTDE) and double TDE (disruption of a stellar binary) scenarios, since they are considered singular events. The pTDE case is unlikely due to inconsistent timing of the flare intervals: the first and second flares occur roughly 110 days apart, while the second and third are separated by about 35 days. Such a dramatic change in timing would require the star’s orbit to be significantly altered. This inconsistency could be explained by a pTDE for the first two flares and accretion disc emission for the third flare.
An alternative interpretation for the observed light curve morphology is the double TDE scenario, followed by a pTDE. In double TDEs, a stellar binary is disrupted by the SMBH (e.g. Mandel & Levin 2015; Mainetti et al. 2016). Mandel & Levin (2015) note that double TDEs could produce a distinct double-peaked signature, where one star is initially disrupted, while the second is captured in an elliptical orbit around the SMBH. This scenario could potentially explain for AT 2022exr, with the first star being fully disrupted and the second one entering an elliptical orbit undergoing partial disruptions, producing the second and third flares. However, Mandel & Levin (2015) found through simulations that the orbital period of the second star can range from 6 months up to several decades, with a median period of roughly 50 years. In the case of AT 2022exr the time interval between the first two flares is only 110 days, shorter than the expected lower limit, making this scenario less likely. Furthermore, Mainetti et al. (2016) modelled light curves of double TDEs using smoothed particle hydrodynamics simulations; however, these simulations did not produce double peaked light curves. This further challenges the double TDE interpretation for AT 2022exr.
4.3.3 AT 2021uvz
Similar to AT 2022exr, AT 2021uvz displays a double-peaked flare, as seen in the upper panel of Fig. 6(e). We observe that both peaks share the same shape, despite their close temporal proximity.
Following our discussion of AT 2022exr’s double-peaked flare, it is highly unlikely that two separate events caused AT 2021uvz’s flares. Under the repeating pTDE hypothesis, we can assume an orbital period of roughly 55 days (the time interval between peaks), which is theoretically possible, with the star being completely disrupted after its second arrival at the pericenter. As for the double TDE scenario, the light curve of AT 2021uvz more closely resembles what one would expect if both stars in the binary system were tidally disrupted almost immediately one after the other. Mandel & Levin (2015) found that 20% of their simulated double TDEs had a time difference in peak luminosities times exceeding 2 months, with individual rise times of less than 1 month. However, the first flare has an observed rise time of 50 days. We note that Mandel & Levin (2015) define the rise time as the duration required for the accretion rate to increase from low values to its peak for each star’s disruption, whereas we refer to the time taken for the light curve to rise from the non-outburst plateau to peak luminosity. Since the origin of optical emission in optical TDEs is not clear, we cannot further assess the plausibility of this scenario.
4.3.4 AT 2020acka
The upper panel of Fig. 7(a) displays the ZTF -band light curve of the AT 2020acka’s flares. The second flare appears more like a bump than a flare, which could result from an accretion disc fluctuation or a sudden influx of excess orbiting material from the disrupted star. The nature of this outburst is not clear from the -band light curve alone. To investigate further, we studied the ZTF -band light curve shown in Fig. 8. Despite the large gap in the -band data, the second flare is more prominent. From the bottom panel of Fig. 7(a) and Fig. 8 we cannot conclusively determine whether the two flares share similar shapes. Consequently, we cannot reliably comment on the nature of the second, bump-like flare.
4.3.5 AT 2019teq
Approximately 300 days prior to the TDE flare, this source exhibited unexpected variability. Although it was included in the Hammerstein et al. (2023) sample, the pre-TDE flare-like feature was not addressed. We present both the ZTF -band and -band light curves in the upper panels of Figs. 7(c) and 7(d) respectively, as artifacts can sometimes affect only one band. Unfortunately there is a gap in the data after the rise of the first flare. However it is evident that the rise time differs significantly from the TDE flare resembling the behaviour observed in AT 2021mhg.
Interestingly, AT 2019teq was detected in X-rays both during the TDE flare period and 3 years later (Yao & Guolo 2022; Ajay et al. 2022). It was also detected in the radio at 6 GHz using the Very Large Array (VLA) (Cendes et al. 2022). Yao & Guolo (2022) and Ajay et al. (2022) suggest that the observed brightening and spectral hardening of the X-rays indicate the formation of a corona. The radio detection further suggests the possible launch of a jet, as only upper limits were obtained in two previous observations.
4.3.6 AT 2019ehz
This TDE was first studied as a part of the van Velzen et al. (2021) sample and was later included in both the Hammerstein et al. (2023) and Yao et al. (2023) samples. However, none of these studies addressed its second flare-like structure. The light curves for both the ZTF -band and -band are presented in the upper panels of Figs. 7(c) and 7(d) respectively. In the bottom panels of these figures, we compare the shapes of the two flares in each filter. Notably, in the -band light curve both flares exhibit a similar shape, consistent with the other repeating TDE-related flares in our sample. However, this is not the case for the -band light curve, where the shapes differ. This behaviour is reminiscent of the flares of AT 2020acka, where the second flare-like bump was more prominent in the -band in than the -band. This behaviour could be due to accretion disc emission but the similarity of the flare shapes in the -band point toward an alternative explanation.
Under the assumption that the second flare is related to the initial TDE, we can investigate its repeating nature. We first assess the likelihood that the two flares resulted from two separate stars being disrupted in the same galaxy. AT 2019ehz was one of the 33 TDEs in the sample of Yao et al. (2023), from which they estimated an optical TDE rate of yr-1 galaxy-1. Following the methodology of Sun et al. (2024), we estimate the probability of a TDE occurring after the initial flare as . Adopting the aforementioned optical TDE rate and the time interval between flares to be days, we obtain . This probability is extremely low. For to exceed , the TDE rate would need to be roughly 1900 times higher, which is far beyond observed values, even for the theoretically enhanced rates in post starburst galaxies. Hence we can confidently exclude the scenario that the two flares were caused by separate events.
In the pTDE case the stellar remnant would need to be in an orbit with a period of 300 days, which is theoretically possible. However, the time separation between the two flares is too large for them to be caused by a full double TDE. The only viable scenario would involve the first star being fully disrupted while the second star is caught in an elliptical orbit around the SMBH, with an orbital period similar to that required for the repeating pTDE interpretation. We note that the first flare’s light curve in Fig. 7(e) displays a knee141414Noticeable change in the slope of the light curve. in its decay, resembling features found in the simulations of Mainetti et al. (2016). In these simulations, such features were only seen in cases where both stars in an equal-mass binary were partially disrupted or in cases with highly unequal-mass binaries.
4.4 Infrared Flares
A small subset (18) of TDEs in our sample display IR flares, along with their optical counterpart, namely : AT 2023ugy, AT 2023cvb, AT 2022upj, AT 2022fpx, AT 2022agi, AT 2022dyt, AT 2022aee, AT 2021uqv, AT 2020afhd, AT 2020ksf, AT 2020mot, AT 2020nov, AT 2020pj, AT 2019qiz, AT 2019dsg, AT 2019azh, AT 2017gge and ASASSN-14li.
In Fig. 9 we show two examples of infrared light curves (lower panels) from the NEOWISE survey (W1 and W2 filters), in addition to optical (middle panels) and X-ray data (upper panels). Only 4 TDEs, namely AT 2023ugy, AT 2022aee, AT 2020mot, AT 2020pj, lacked an X-ray counterpart (with only upper limits available). The rest of the light curves can be plotted using the accompanying app that is available on the GitHub page.
4.5 UV Emission
Following the process described in Sect. 3.3, we retrieved UVOT photometry for 118 TDEs. An example light curve (AT 2020nov) is displayed in Fig. 10. In total, 111 TDEs have data for all 6 wavebands, while 95 have at least 5 observations for each of the three UV bands. These data can be accessed through the GitHub page, under the UVOT folder in the photometry section. Additionally, these light curves can be plotted using the accompanying app, also available on the GitHub page.
4.6 X-ray outbursts
We also investigated TDEs observed in X-rays (see Sect. 3.2). In our main sample, 122 sources have either detections or 3 upper limits in the X-rays. Of these, 45 TDEs have at least 1 detection, while 26 have at least 5 detections. Example light curves are shown in the upper panels of Fig. 9. All available X-ray data can be found in the accompanying GitHub repository and plotted using the provided app. We note that we analyse X-ray observations individually, meaning that stacking observations will lead to more detections. We also consider as detections only the observations with signal-to-noise (SNR) ratio equal to or greater than 3. The SNR of the observations is also included in the GitHub files of each source, so the users can set the limit to the threshold they want.
5 Discussion
5.1 Statistics of the sample
In Sect. 4.1, we showed that the distributions of the duration, and , as well as their ratio, are consistent with a log-normal distribution. Log-normal distributions are usually encountered in systems driven by multiplicative processes or exponentially growing or decaying dynamics. They also emerge when the observed data result from a combination of independent processes with varying initial conditions. This means that the diversity in stellar properties and BH environments in TDEs could be responsible for the log-normal distributions.
The log-normal behaviour of the redshift distribution could result from observational selection effects rather than from some intrinsic property of TDEs. We expect that the distribution of redshifts will change with the addition of data from upcoming surveys such as the Legacy Survey of Space and Time151515https://www.lsst.org/about (LSST; Ivezic et al. 2008).
5.2 Spectral classes
As shown in Fig. 5, the majority of TDEs fall into the TDE-H+He category. As predicted by Syer & Ulmer (1999), the disruption of main-sequence (MS) stars by SMBHs of mass would be the easiest for modern-day telescopes to observe. MS stars consist of both hydrogen and helium, hence the observed spectral signatures. Moreover, MS stars are abundant in galactic centres due to their long-lived nature and have been found to dominate TDE rates in past studies (e.g. Kochanek 2016). The second most common spectral class in our sample is TDE-H, followed by TDE-He and TDE-featureless classes. The statistical study of Nicholl et al. (2022) uncovered that for the TDE-H sources in their sample, their impact parameter (b; i.e. high b almost full disruption, small b partial disruption) was by far the smallest compared to TDE-H+He and TDE-He, corresponding to partial disruptions. The outer layers of MS stars are mainly hydrogen, so the partial disruption of the outer layers of MS stars could explain the observed TDE-H population. Additionally, hydrogen-rich young MS stars are typically rare, hence the contribution of their disruption to the population should be small. On the other hand, the TDE-He class is significantly more scarcely observed, with only 16 events (18 if we consider the different epoch classifications from Sect. 4.2). This could be explained by the disruption of helium-rich stellar cores, as suggested by Gezari et al. (2012) for PS1-10jh. These stellar remnants are rare, theorized to be originating from stellar binary systems (e.g. Paczyński 1967; Iben & Tutukov 1985; Götberg et al. 2019), which could explain their small observed number in TDEs.
Regarding the TDE-featureless class, we have collected data for 13 such events. Hammerstein et al. (2023) found that the 4 TDE-featureless sources in their sample exhibited higher peak bolometric luminosities, hotter temperatures and larger radii than those in the TDE-H and TDE-H+He classes. Moreover, their TDE-featureless sample favoured redder and more massive host galaxies compared to other classes. This trend can be further tested using the TDEs in our catalogue. Featureless TDE spectra can arise for several reasons. One possibility is that the line-emitting regions are obscured, depending on our viewing angle. Another explanation is that dense, optically thick outflows reprocess and thermalize high-energy photons, effectively ”washing out” sharp emission lines. Additionally, a strong, hot continuum can overpower weaker lines, making them difficult to detect.
To analyse the variability properties of different spectral classes, we applied the same statistical approach used in Sect. 4.1 to the TDEs for which we could reliably measure flare variability. We then utilized a K-S test to assess whether the distributions of different spectral classes are consistent with that of the full sample. We obtained p-values ranging from 0.15 to 0.97 for the rise times, decay times and durations. Although the sample sizes for TDE-H, TDE-He and TDE-featureless classes are small, we find that the time scale distributions of all spectral classes are drawn from the same parent population.
Additionally, we investigated the spectral classes of TDEs in different subsamples showing either IR emission, repeating flares or X-ray emission. For the IR sample we find that 12 are TDE-H+He, 5 are TDE-H and 1 is TDE-He. For the repeating flare sample we find that 8 are TDE-H+He, 2 are TDE-H and 2 are TDE-featureless. Finally, for the X-ray detection sample we find that 4 are TDE-He, 26 are TDE-H+He, 9 are TDE-H and 3 are TDE-featureless. We estimated the probability that the spectral classes of the different subsamples are randomly drawn from the full sample using:
where is the binomial coefficient. Here, the numerator represents the number of ways to compose a sample with the different spectral classes in each subsample while the denominator accounts for all possible selections from the full set of 134 TDEs. The small p-values in all three cases indicate the spectral composition of these subsamples is unlikely to be random. This could indicate that the spectral characteristics of TDEs are related to their overall multi-wavelength behaviour.
5.3 Similarities and specific cases in the repeating flare sample
From our main catalogue, 12 TDEs exhibit more than one flare. While half of these TDEs have been extensively studied in previous works, 3 of the remaining 6 sources (AT 2020acka, AT 2019teq, and AT 2019ehz) have appeared in past sample studies (van Velzen et al. 2021; Hammerstein et al. 2023; Yao et al. 2023) without their repeating nature being highlighted. Only recently were AT 2020acka’s and AT 2019ehz’s re-brightening features referenced by Guo et al. (2025) and Zhong (2025). Unfortunately, due to a lack of spectroscopic data, we cannot confirm the TDE nature of the flares in the latter half of our repeating flare sample. However, by comparing the rise times, decay times and durations of both flares we find that the timescales for AT 2022dbl, AT 2022agi, AT 2020vdq, AT 2019ehz, and AT 2019azh, are consistent within uncertainty with one another (see Appendix D). We are unable to compute the rise time of the second flare of AT 2021mhg and the decay time of the second flare of AT 2019teq due to substantial gaps in the respective light curves. Additionally, for AT2022exr and AT2021uvz, which are the double peak flares, as well as AT2020acka, we cannot evaluate the shape of their flares since they are overlapping. From a visual inspection, the rise time of the second flare of AT 2021mhg is clearly different and has been identified as a Type Ia SN. Moreover, the flares in AT 2020acka and AT 2019teq, possibly originating from accretion disc emission and an unknown transient respectively, also display dissimilar shapes. The similarity of the repeating flares has also been previously noted by Makrygianni et al. (2025) for AT 2022dbl and Hinkle et al. (2021) for AT 2019azh.
If the additional flares arise from pTDEs, one would expect their shapes to be similar since they would result from the disruption of the same star. However, the observed diversity in flare shapes suggests that underlying mechanisms may be at play: some flares could indeed be repeating pTDEs, while others may result from distinct events such as double TDEs or unrelated transient phenomena. For instance, the 18-year interval between flares in AT 2024pvu makes it unlikely that both events originate from the partial disruption of the same star. In this context, the overall shape of a flare may be influenced primarily by the properties of the SMBH rather than by the characteristics of the disrupted star. Moreover, Liu et al. (2025) found in their simulations that a sun-like star would produce multiple flares of increasing luminosity, consistent with AT 2020vdq, whose second flare is more luminous than the first. However, this picture is not consistent with AT 2022dbl, where the opposite behaviour is observed.
Furthermore, 8 out of the 12 sources in our repeating flare sample exhibit X-ray detections with the exception of AT 2024pvu, AT 2022dbl, AT 2021mhg, and AT 2020acka, for which we obtained only upper limits. We observe that in some TDEs (AT 2022exr,,AT 2020vdq, AT 2019ehz, AT 2019teq, and AT 2018fyk), the X-ray emission follows the optical flare with only a short delay. In contrast, others (AT 2021uvz, AT 2022agi, and AT 2019azh) display significantly longer delays. Notably, for AT 2022exr and AT 2020vdq, X-rays are detected only after the second flare, while for AT 2018fyk they occur nearly simultaneously with the optical outbursts.
5.4 IR TDE candidate samples
In addition to the confirmed TDEs presented in Sect. 4.4, several studies have identified samples of infrared-selected TDE candidates. For example, Masterson et al. (2024) identified 18 candidates in nearby galactic nuclei (within 200 Mpc) based on distinctive mid-infrared light curve features. They divided these into two subsamples—a set of six candidates where the variability might be linked to an underlying AGN, and a set of 12 more robust candidates (including WTP14adbjsh, initially discussed in Panagiotou et al. 2023). Similarly, Jiang et al. (2021) examined IR flares in SDSS galaxies, identifying 137 mid-infrared outbursts, of which two have been proposed as TDE candidates. TDECat provides a sample of optical TDEs with an IR counterpart which can be used for comparison with IR-selected TDE candidate samples to investigate if a potential connection exists.
5.5 Correlation between X-ray and Infrared emission?
As shown in Sect. 4.4, most TDEs exhibiting IR flares also display X-ray detections. In fact, 14 out of the 18 IR-flaring TDEs have been detected in X-rays. Moreover, in the sample of infrared-selected TDEs from the eROSITA sky presented by Masterson et al. (2024), 3 of the 8 objects (all from their gold sample) were detected in the X-rays at multiple epochs. These findings suggest a potential underlying correlation between IR and X-ray emission.
Notably, both confirmed TDEs associated with astrophysical neutrinos, AT 2019dsg (Stein et al. 2021a) and AT 2017gge (Li et al. 2024), are included in our IR flare sample. It has been proposed that neutrino production in these events could be linked to radio emission (Stein et al. 2021a), infrared emission (Yuan et al. 2024), or X-ray emission (Li et al. 2024). However, further study is needed to clarify the association between neutrinos and TDEs, as well as the particle acceleration mechanisms that might lead to neutrino production.
We also investigated whether any sources exhibited delayed IR or X-ray emission relative to their optical flares. Our analysis reveals that in two TDEs (AT 2021uqv, AT 2019dsg) only the IR emission is delayed; in another five cases (AT 2022fpx, AT 2020ksf, AT 2020nov, AT 2020afhd, AT 2019azh), only the X-rays are delayed; and in five cases, both IR and X-ray emissions are delayed, with IR flares generally lasting longer, and the X-ray emission being more delayed. In contrast, for AT 2022dyt and ASASSN-14li, the X-ray and IR emission occur almost simultaneously with the optical flare.
Further similarities emerge within our sample. For example, AT 2022fpx, extensively studied by Koljonen et al. (2024), displays a light curve similar to that of AT 2020nov, where the X-ray emission coincides with a bump in the optical light curve. Furthermore, all coronal line–emitting TDEs, namely, AT 2022fpx (Koljonen et al. 2024), AT 2022upj (Newsome et al. 2022, 2024), AT 2019qiz (Short et al. 2023), and AT 2017gge (Onori et al. 2022), exhibit both X-ray and IR emission.
We can estimate the statistical significance of the correlation between IR and X-ray emission in our sample of optical TDEs, by considering a binomial calculation. The chance that a TDE has X-ray emission can be calculated as the detection fraction of the X-rays in our optical TDE sample (i.e. p=45/134=0.336). We can then test the null hypothesis (i.e. that each of the IR-flaring TDEs has an independent chance p=0.336 of X-ray detection), by computing the probability of seeing X-ray detections by chance
Since , we reject the null hypothesis, meaning that the observed correlation of X-ray and IR emission in our IR-emitting optical TDE sample is likely not by chance. We can also convert the probability to sigma levels for a one-tailed normal distribution, which gives a 3.6 significance.
5.6 X-ray TDE candidate samples
Previous works have studied the X-ray TDE candidate population (e.g. Auchettl et al. 2017; Sazonov et al. 2021; Khorunzhev et al. 2022). However, there is still uncertainty on the properties that distinguish X-ray TDEs from active galactic nuclei (AGN) activity. The TDEs in our X-ray sample, which also have an optical counterpart (except from the 3 jetted TDEs), provide a valuable opportunity to refine the classification criteria for X-ray TDEs that lack detectable optical emission. Moreover, the substantially larger number of X-ray bright optical TDEs now available enables more comprehensive comparative studies with the AGN population (e.g. Auchettl et al. 2018), which may ultimately help elucidate the intrinsic differences between these phenomena.
6 Conclusions
We have compiled a catalogue of 134 confirmed TDEs up until the end of 2024. We collected multi-wavelength photometry (X-ray, UV, optical, and infrared) along with publicly available spectra. The complete dataset is accessible via a dedicated GitHub repository and a Python application. Additionally, we created a list of strong TDE candidates.
We analysed the sources in our main sample by investigating their statistical properties, spectral classes, and the presence of repeating flares, IR emission, and X-ray outbursts. Our key findings are as follows:
1.
By implementing a custom Bayesian block algorithm, we estimated the duration, rise time, decay time, and their ratio for the optical flares of 101 TDEs in our sample. A total of 103 flares were analysed, including all flares from repeating pTDEs (AT 2022dbl and AT 2020vdq). We find that the distributions of the ratio, durations, rise times, and decay times are well described by a log-normal distribution.
2.
The best-fit line for vs is given by , where and , with 95% confidence intervals of 0.8374 to 1.1227 and -0.6859 to -0.0168 for and respectively. This result indicates that the majority of the TDE population is confined within the scatter of this best-fit line. We also find a positive correlation between and . These results, in combination with the log-normal distributions of the timescales, can be used to create artificial light curves for future statistical studies.
3.
Our spectral analysis reveals that the majority of TDEs belong to the TDE-H+He class, followed by the TDE-H class, with the TDE-He and TDE-featureless occurring at slightly lower frequencies. This distribution is consistent with expectations based on the predominance of main sequence stars in galactic centres. Although we computed timescale statistics for the different spectral classes, we did not identify any definite trends.
4.
We examined whether specific spectral classes tend to exhibit repeating flares, infrared emission, or X-ray outbursts. Our analysis reveals that the spectral class distribution within these TDE subsamples is unlikely to have arisen by random chance.
5.
We identified 4 new TDEs that show secondary flares in their optical light curves, namely AT 2024pvu, AT 2022exr, AT 2021uvz, and AT 2019teq. We further studied recently referenced TDEs displaying re-brightenings: namely AT 2020acka, and AT 2019ehz. We also compared the shape of the multiple flares for each TDE and found that their shapes are generally similar, excluding AT 2021mhg, AT 2020acka and AT 2019teq, which exhibit distinct differences.
6.
We observe a potential correlation between infrared and X-ray emission; most TDEs (14 out of 18) displaying infrared flares also exhibit X-ray detections. We plan to explore this finding further in a future study. Moreover, all reported coronal line emitting TDEs were found to exhibit both X-ray and IR emission.
This comprehensive (mostly optical) TDE catalogue not only provides a robust dataset for statistical studies and machine learning applications, but also paves the way for future population studies aimed at understanding the characteristics of still poorly understood TDEs. As we enter into the data ”rich” era, especially with forthcoming surveys like the LSST, which are expected to capture tens of TDEs each night, this work serves as an essential tool for preparing for the challenges and opportunities ahead.
Acknowledgements.
The authors thank the anonymous referee for comments and suggestions that helped improve this work. D.A.L., A.P., and I.L. were funded by the European Union ERC-2022-STG - BOOTES - 101076343. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. N.G. acknowledge support by a grant from the Simons Foundation (00001470, NG). J.D., D.M., and Z.T. took part in this research under the auspices of the Science Internship Program at the University of California Santa Cruz. N.G. is grateful to Caitlyn Nojiri and Hannah Dykaar for their help. BHTOM has been based on the open-source TOM Toolkit by LCO and has been developed with funding from the European Union’s Horizon 2020 and Horizon Europe research and innovation programmes under grant agreements No 101004719 (OPTICON-RadioNet Pilot, ORP) and No 101131928 (ACME). We acknowledge ESA Gaia, DPAC and the Photometric Science Alerts Team (http://gsaweb.ast.cam.ac.uk/alerts). This work has made use of data from the Asteroid Terrestrial-impact Last Alert System (ATLAS) project. The Asteroid Terrestrial-impact Last Alert System (ATLAS) project is primarily funded to search for near earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogs from the survey area. This work was partially funded by Kepler/K2 grant J1944/80NSSC19K0112 and HST GO-15889, and STFC grants ST/T000198/1 and ST/S006109/1. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen’s University Belfast, the Space Telescope Science Institute, the South African Astronomical Observatory, and The Millennium Institute of Astrophysics (MAS), Chile. The Pan-STARRS1 Surveys (PS1) and the PS1 public science archive have been made possible through contributions by the Institute for Astronomy, the University of Hawaii, the Pan-STARRS Project Office, the Max-Planck Society and its participating institutes, the Max Planck Institute for Astronomy, Heidelberg and the Max Planck Institute for Extraterrestrial Physics, Garching, The Johns Hopkins University, Durham University, the University of Edinburgh, the Queen’s University Belfast, the Harvard-Smithsonian Center for Astrophysics, the Las Cumbres Observatory Global Telescope Network Incorporated, the National Central University of Taiwan, the Space Telescope Science Institute, the National Aeronautics and Space Administration under Grant No. NNX08AR22G issued through the Planetary Science Division of the NASA Science Mission Directorate, the National Science Foundation Grant No. AST-1238877, the University of Maryland, Eotvos Lorand University (ELTE), the Los Alamos National Laboratory, and the Gordon and Betty Moore Foundation. The CSS survey is funded by the National Aeronautics and Space Administration under Grant No. NNG05GF22G issued through the Science Mission Directorate Near-Earth Objects Observations Program. The CRTS survey is supported by the U.S. National Science Foundation under grants AST-0909182. This publication makes use of data products from NEOWISE, which is a project of the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration. This publication makes use of data products from the Wide-field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration. This work utilises Matplotlib (Hunter 2007), NumPy (Harris et al. 2020) and SciPy (Virtanen et al. 2020).References
- Ajay et al. (2022) Ajay, Y., Pasham, D., & Guolo, M. 2022, The Astronomer’s Telegram, 15724, 1
- Alexander et al. (2016) Alexander, K. D., Berger, E., Guillochon, J., Zauderer, B. A., & Williams, P. K. G. 2016, ApJ, 819, L25
- Alexander et al. (2021) Alexander, K. D., Velzen, S. V., Miller-Jones, J., et al. 2021, Transient Name Server AstroNote, 24, 1
- Andreoni et al. (2022) Andreoni, I., Coughlin, M. W., Perley, D. A., et al. 2022, Nature, 612, 430
- Arcavi et al. (2014) Arcavi, I., Gal-Yam, A., Sullivan, M., et al. 2014, ApJ, 793, 38
- Auchettl et al. (2017) Auchettl, K., Guillochon, J., & Ramirez-Ruiz, E. 2017, ApJ, 838, 149
- Auchettl et al. (2018) Auchettl, K., Ramirez-Ruiz, E., & Guillochon, J. 2018, ApJ, 852, 37
- Bade et al. (1996) Bade, N., Komossa, S., & Dahlem, M. 1996, A&A, 309, L35
- Bellm et al. (2019) Bellm, E. C., Kulkarni, S. R., Graham, M. J., et al. 2019, PASP, 131, 018002
- Blagorodnova et al. (2019) Blagorodnova, N., Cenko, S. B., Kulkarni, S. R., et al. 2019, ApJ, 873, 92
- Blanton & Roweis (2007) Blanton, M. R. & Roweis, S. 2007, AJ, 133, 734
- Bloom et al. (2011) Bloom, J. S., Giannios, D., Metzger, B. D., et al. 2011, Science, 333, 203
- Brimacombe et al. (2015) Brimacombe, J., Brown, J. S., Holoien, T. W. S., et al. 2015, The Astronomer’s Telegram, 7910, 1
- Brown et al. (2015) Brown, G. C., Levan, A. J., Stanway, E. R., et al. 2015, MNRAS, 452, 4297
- Burrows et al. (2005) Burrows, D. N., Hill, J. E., Nousek, J. A., et al. 2005, Space Sci. Rev., 120, 165
- Burrows et al. (2011) Burrows, D. N., Kennea, J. A., Ghisellini, G., et al. 2011, Nature, 476, 421
- Capalbi et al. (2005) Capalbi, M., Perri, M., Saija, B., Tamburelli, F., & Angelini, L. 2005, Version, 1, 28
- Cendes et al. (2022) Cendes, Y., Alexander, K., Berger, E., et al. 2022, The Astronomer’s Telegram, 15728, 1
- Cenko et al. (2016) Cenko, S. B., Cucchiara, A., Roth, N., et al. 2016, ApJ, 818, L32
- Cenko et al. (2012) Cenko, S. B., Krimm, H. A., Horesh, A., et al. 2012, ApJ, 753, 77
- Chambers et al. (2016) Chambers, K. C., Magnier, E. A., Metcalfe, N., et al. 2016, arXiv e-prints, arXiv:1612.05560
- Charalampopoulos et al. (2022) Charalampopoulos, P., Leloudas, G., Malesani, D. B., et al. 2022, A&A, 659, A34
- Charalampopoulos et al. (2023) Charalampopoulos, P., Leloudas, G., Pursiainen, M., & Kotak, R. 2023, Transient Name Server AstroNote, 115, 1
- D’Elia et al. (2013) D’Elia, V., Perri, M., Puccetti, S., et al. 2013, A&A, 551, A142
- Drake et al. (2009) Drake, A. J., Djorgovski, S. G., Mahabal, A., et al. 2009, ApJ, 696, 870
- Fitzpatrick & Massa (2007) Fitzpatrick, E. L. & Massa, D. 2007, ApJ, 663, 320
- Flewelling et al. (2020) Flewelling, H. A., Magnier, E. A., Chambers, K. C., et al. 2020, ApJS, 251, 7
- Freeman et al. (2001) Freeman, P., Doe, S., & Siemiginowska, A. 2001, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 4477, Astronomical Data Analysis, ed. J.-L. Starck & F. D. Murtagh, 76–87
- Fruscione et al. (2006) Fruscione, A., McDowell, J. C., Allen, G. E., et al. 2006, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 6270, Observatory Operations: Strategies, Processes, and Systems, ed. D. R. Silva & R. E. Doxsey, 62701V
- Gehrels (1986) Gehrels, N. 1986, ApJ, 303, 336
- Gezari et al. (2012) Gezari, S., Chornock, R., Rest, A., et al. 2012, Nature, 485, 217
- Gilfanov et al. (2020) Gilfanov, M., Sazonov, S., Sunyaev, R., et al. 2020, The Astronomer’s Telegram, 14246, 1
- Götberg et al. (2019) Götberg, Y., de Mink, S. E., Groh, J. H., Leitherer, C., & Norman, C. 2019, A&A, 629, A134
- Graham et al. (2019) Graham, M. J., Kulkarni, S. R., Bellm, E. C., et al. 2019, PASP, 131, 078001
- Grupe et al. (1999) Grupe, D., Thomas, H. C., & Leighly, K. M. 1999, A&A, 350, L31
- Guo et al. (2025) Guo, H., Sun, J., Li, S., et al. 2025, ApJ, 979, 235
- Hammerstein et al. (2023) Hammerstein, E., van Velzen, S., Gezari, S., et al. 2023, ApJ, 942, 9
- Harris et al. (2020) Harris, C. R., Millman, K. J., Van Der Walt, S. J., et al. 2020, Nature, 585, 357
- Hart et al. (2023) Hart, K., Shappee, B. J., Hey, D., et al. 2023, arXiv e-prints, arXiv:2304.03791
- Heinze et al. (2018) Heinze, A. N., Tonry, J. L., Denneau, L., et al. 2018, AJ, 156, 241
- Hills (1975) Hills, J. G. 1975, Nature, 254, 295
- Hinkle et al. (2021) Hinkle, J. T., Holoien, T. W. S., Auchettl, K., et al. 2021, MNRAS, 500, 1673
- Ho et al. (2025) Ho, A. Y. Q., Yao, Y., Matsumoto, T., et al. 2025, arXiv e-prints, arXiv:2502.07885
- Hodgkin et al. (2021) Hodgkin, S. T., Harrison, D. L., Breedt, E., et al. 2021, A&A, 652, A76
- Holoien et al. (2016a) Holoien, T. W. S., Kochanek, C. S., Prieto, J. L., et al. 2016a, MNRAS, 463, 3813
- Holoien et al. (2016b) Holoien, T. W. S., Kochanek, C. S., Prieto, J. L., et al. 2016b, MNRAS, 455, 2918
- Holoien et al. (2014) Holoien, T. W. S., Prieto, J. L., Bersier, D., et al. 2014, MNRAS, 445, 3263
- Hung et al. (2017) Hung, T., Gezari, S., Blagorodnova, N., et al. 2017, ApJ, 842, 29
- Hunter (2007) Hunter, J. D. 2007, Computing in Science and Engineering, 9, 90
- Iben & Tutukov (1985) Iben, Jr., I. & Tutukov, A. V. 1985, ApJS, 58, 661
- Inserra et al. (2012) Inserra, C., Smartt, S. J., Fraser, M., et al. 2012, The Astronomer’s Telegram, 4329, 1
- Ivezic et al. (2008) Ivezic, Z., Axelrod, T., Brandt, W. N., et al. 2008, Serbian Astronomical Journal, 176, 1
- Jiang et al. (2016) Jiang, N., Dou, L., Wang, T., et al. 2016, ApJ, 828, L14
- Jiang et al. (2021) Jiang, N., Wang, T., Dou, L., et al. 2021, ApJS, 252, 32
- Jose et al. (2014) Jose, J., Guo, Z., Long, F., et al. 2014, The Astronomer’s Telegram, 6777, 1
- Kalberla et al. (2005) Kalberla, P. M. W., Burton, W. B., Hartmann, D., et al. 2005, A&A, 440, 775
- Khorunzhev et al. (2022) Khorunzhev, G. A., Sazonov, S. Y., Medvedev, P. S., et al. 2022, Astronomy Letters, 48, 767
- Kochanek (2016) Kochanek, C. S. 2016, MNRAS, 461, 371
- Kochanek et al. (2017) Kochanek, C. S., Shappee, B. J., Stanek, K. Z., et al. 2017, PASP, 129, 104502
- Koljonen et al. (2024) Koljonen, K. I. I., Liodakis, I., Lindfors, E., et al. 2024, MNRAS, 532, 112
- Krimm et al. (2011) Krimm, H. A., Kennea, J. A., Holland, S. T., et al. 2011, The Astronomer’s Telegram, 3384, 1
- Kumar et al. (2024) Kumar, H., Berger, E., Hiramatsu, D., et al. 2024, ApJ, 974, L36
- Law et al. (2009) Law, N. M., Kulkarni, S. R., Dekany, R. G., et al. 2009, PASP, 121, 1395
- Li et al. (2024) Li, R.-L., Yuan, C., He, H.-N., et al. 2024, arXiv e-prints, arXiv:2411.06440
- Liddle (2007) Liddle, A. R. 2007, MNRAS, 377, L74
- Lin et al. (2024) Lin, Z., Jiang, N., Wang, T., et al. 2024, ApJ, 971, L26
- Liu et al. (2025) Liu, C., Yarza, R., & Ramirez-Ruiz, E. 2025, ApJ, 979, 40
- Liu et al. (2017) Liu, T., Tozzi, P., Wang, J.-X., et al. 2017, ApJS, 232, 8
- Magnier et al. (2020a) Magnier, E. A., Chambers, K. C., Flewelling, H. A., et al. 2020a, ApJS, 251, 3
- Magnier et al. (2020b) Magnier, E. A., Schlafly, E. F., Finkbeiner, D. P., et al. 2020b, ApJS, 251, 6
- Magnier et al. (2020c) Magnier, E. A., Sweeney, W. E., Chambers, K. C., et al. 2020c, ApJS, 251, 5
- Mainetti et al. (2016) Mainetti, D., Lupi, A., Campana, S., & Colpi, M. 2016, MNRAS, 457, 2516
- Mainzer et al. (2014) Mainzer, A., Bauer, J., Cutri, R. M., et al. 2014, ApJ, 792, 30
- Mainzer et al. (2011) Mainzer, A., Bauer, J., Grav, T., et al. 2011, ApJ, 731, 53
- Makrygianni et al. (2025) Makrygianni, L., Arcavi, I., Newsome, M., et al. 2025, arXiv e-prints, arXiv:2505.16867
- Mandel & Levin (2015) Mandel, I. & Levin, Y. 2015, ApJ, 805, L4
- Marchesi et al. (2016) Marchesi, S., Lanzuisi, G., Civano, F., et al. 2016, ApJ, 830, 100
- Masci et al. (2019) Masci, F. J., Laher, R. R., Rusholme, B., et al. 2019, PASP, 131, 018003
- Masterson et al. (2024) Masterson, M., De, K., Panagiotou, C., et al. 2024, ApJ, 961, 211
- Merloni et al. (2024) Merloni, A., Lamer, G., Liu, T., et al. 2024, A&A, 682, A34
- Miller et al. (2015) Miller, J. M., Kaastra, J. S., Miller, M. C., et al. 2015, Nature, 526, 542
- Monet et al. (2003) Monet, D. G., Levine, S. E., Canzian, B., et al. 2003, AJ, 125, 984
- Moretti et al. (2005) Moretti, A., Campana, S., Mineo, T., et al. 2005, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 5898, UV, X-Ray, and Gamma-Ray Space Instrumentation for Astronomy XIV, ed. O. H. W. Siegmund, 360–368
- Munoz-Arancibia et al. (2023) Munoz-Arancibia, A., Hernandez-Garcia, L., Bauer, F. E., et al. 2023, Transient Name Server AstroNote, 138, 1
- Nevalainen et al. (2005) Nevalainen, J., Markevitch, M., & Lumb, D. 2005, ApJ, 629, 172
- Newsome et al. (2022) Newsome, M., Arcavi, I., Dgany, Y., & Pellegrino, C. 2022, Transient Name Server AstroNote, 236, 1
- Newsome et al. (2024) Newsome, M., Arcavi, I., Howell, D. A., et al. 2024, ApJ, 977, 258
- Nicholl et al. (2019) Nicholl, M., Blanchard, P. K., Berger, E., et al. 2019, MNRAS, 488, 1878
- Nicholl et al. (2022) Nicholl, M., Lanning, D., Ramsden, P., et al. 2022, MNRAS, 515, 5604
- Oke & Gunn (1983) Oke, J. B. & Gunn, J. E. 1983, ApJ, 266, 713
- Onori et al. (2022) Onori, F., Cannizzaro, G., Jonker, P. G., et al. 2022, MNRAS, 517, 76
- Paczyński (1967) Paczyński, B. 1967, Acta Astron., 17, 355
- Paggi et al. (2013) Paggi, A., Massaro, F., D’Abrusco, R., et al. 2013, ApJS, 209, 9
- Panagiotou et al. (2023) Panagiotou, C., De, K., Masterson, M., et al. 2023, ApJ, 948, L5
- Pasham et al. (2017) Pasham, D. R., Cenko, S. B., Sadowski, A., et al. 2017, ApJ, 837, L30
- Pasham et al. (2023) Pasham, D. R., Lucchini, M., Laskar, T., et al. 2023, Nature Astronomy, 7, 88
- Perri et al. (2007) Perri, M., Maselli, A., Giommi, P., et al. 2007, A&A, 462, 889
- Poidevin et al. (2023) Poidevin, F., Pérez-Fournon, I., Geier, S., et al. 2023, Transient Name Server Classification Report, 2023-3050, 1
- Prieto et al. (2014) Prieto, J. L., Bersier, D., Holoien, T. W. S., et al. 2014, The Astronomer’s Telegram, 5831, 1
- Puccetti et al. (2011) Puccetti, S., Capalbi, M., Giommi, P., et al. 2011, A&A, 528, A122
- Rau et al. (2009) Rau, A., Kulkarni, S. R., Law, N. M., et al. 2009, PASP, 121, 1334
- Rees (1988) Rees, M. J. 1988, Nature, 333, 523
- Reis et al. (2012) Reis, R. C., Reynolds, M. T., Miller, J. M., Miller, B. P., & Hadjiyska, E. 2012, The Astronomer’s Telegram, 4346, 1
- Ricci et al. (2017) Ricci, C., Trakhtenbrot, B., Koss, M. J., et al. 2017, ApJS, 233, 17
- Roming et al. (2005) Roming, P. W. A., Kennedy, T. E., Mason, K. O., et al. 2005, Space Sci. Rev., 120, 95
- Saxton et al. (2020) Saxton, R., Komossa, S., Auchettl, K., & Jonker, P. G. 2020, Space Sci. Rev., 216, 85
- Sazonov et al. (2021) Sazonov, S., Gilfanov, M., Medvedev, P., et al. 2021, MNRAS, 508, 3820
- Scargle (1998) Scargle, J. D. 1998, ApJ, 504, 405
- Scargle et al. (2013) Scargle, J. D., Norris, J. P., Jackson, B., & Chiang, J. 2013, ApJ, 764, 167
- Schlafly & Finkbeiner (2011) Schlafly, E. F. & Finkbeiner, D. P. 2011, ApJ, 737, 103
- Shappee et al. (2014) Shappee, B. J., Prieto, J. L., Grupe, D., et al. 2014, ApJ, 788, 48
- Shingles et al. (2021) Shingles, L., Smith, K. W., Young, D. R., et al. 2021, Transient Name Server AstroNote, 7, 1
- Short et al. (2023) Short, P., Lawrence, A., Nicholl, M., et al. 2023, MNRAS, 525, 1568
- Smartt et al. (2012) Smartt, S. J., Inserra, C., Fraser, M., et al. 2012, The Astronomer’s Telegram, 4299, 1
- Smith et al. (2020) Smith, K. W., Smartt, S. J., Young, D. R., et al. 2020, PASP, 132, 085002
- Somalwar et al. (2023a) Somalwar, J. J., Ravi, V., Dong, D. Z., et al. 2023a, arXiv e-prints, arXiv:2310.03791
- Somalwar et al. (2023b) Somalwar, J. J., Ravi, V., Yao, Y., et al. 2023b, arXiv e-prints, arXiv:2310.03782
- Stein et al. (2021a) Stein, R., van Velzen, S., Kowalski, M., et al. 2021a, Nature Astronomy, 5, 510
- Stein et al. (2021b) Stein, R., van Velzen, S., Kowalski, M., et al. 2021b, Nature Astronomy, 5, 510
- Stokes et al. (2000) Stokes, G. H., Evans, J. B., Viggh, H. E. M., Shelly, F. C., & Pearce, E. C. 2000, Icarus, 148, 21
- Subrayan et al. (2023) Subrayan, B. M., Milisavljevic, D., Chornock, R., et al. 2023, ApJ, 948, L19
- Sun et al. (2024) Sun, L., Jiang, N., Dou, L., et al. 2024, Recurring tidal disruption events a decade apart in IRAS F01004-2237
- Syer & Ulmer (1999) Syer, D. & Ulmer, A. 1999, MNRAS, 306, 35
- Tadhunter et al. (2017) Tadhunter, C., Spence, R., Rose, M., Mullaney, J., & Crowther, P. 2017, Nature Astronomy, 1, 0061
- Tonry et al. (2018) Tonry, J. L., Denneau, L., Heinze, A. N., et al. 2018, PASP, 130, 064505
- Tozzi et al. (2006) Tozzi, P., Gilli, R., Mainieri, V., et al. 2006, A&A, 451, 457
- van Velzen et al. (2016) van Velzen, S., Anderson, G. E., Stone, N. C., et al. 2016, Science, 351, 62
- van Velzen et al. (2021) van Velzen, S., Gezari, S., Hammerstein, E., et al. 2021, ApJ, 908, 4
- Virtanen et al. (2020) Virtanen, P., Gommers, R., Oliphant, T. E., et al. 2020, Nature Methods, 17, 261
- Wang et al. (2022) Wang, Y., Jiang, N., Wang, T., et al. 2022, The Astrophysical Journal Letters, 930, L4
- Waters et al. (2020) Waters, C. Z., Magnier, E. A., Price, P. A., et al. 2020, ApJS, 251, 4
- Webb et al. (2020) Webb, N. A., Coriat, M., Traulsen, I., et al. 2020, A&A, 641, A136
- Wevers et al. (2023) Wevers, T., Coughlin, E. R., Pasham, D. R., et al. 2023, ApJ, 942, L33
- Wevers et al. (2019) Wevers, T., Pasham, D. R., van Velzen, S., et al. 2019, MNRAS, 488, 4816
- Wright et al. (2010) Wright, E. L., Eisenhardt, P. R. M., Mainzer, A. K., et al. 2010, AJ, 140, 1868
- Wyrzykowski et al. (2017) Wyrzykowski, Ł., Zieliński, M., Kostrzewa-Rutkowska, Z., et al. 2017, MNRAS, 465, L114
- Yao & Guolo (2022) Yao, Y. & Guolo, M. 2022, The Astronomer’s Telegram, 15657, 1
- Yao et al. (2023) Yao, Y., Ravi, V., Gezari, S., et al. 2023, ApJ, 955, L6
- Yaron & Gal-Yam (2012) Yaron, O. & Gal-Yam, A. 2012, PASP, 124, 668
- York et al. (2000) York, D. G., Adelman, J., Anderson, Jr., J. E., et al. 2000, AJ, 120, 1579
- Yuan et al. (2024) Yuan, C., Winter, W., & Lunardini, C. 2024, ApJ, 969, 136
- Zauderer et al. (2011) Zauderer, B. A., Berger, E., Soderberg, A. M., et al. 2011, Nature, 476, 425
- Zhong (2025) Zhong, S. 2025, ApJ, 983, 131
Appendix A Published TDEs
AT 2023vto: This object was initially classified (and still is) as a superluminous super nova Type II (SLSN II) by Poidevin et al. (2023). This classification was based on faint emission that was detected in optical spectra obtained on 2023-11-21 22:30:53 (UTC). However, Kumar et al. (2024) classified it as a TDE, after identifying the broad emission centred at (see their Fig. 4) to be a broad, blueshifted He II emission line (TDE-He; see Sect. 4.2).
AT 2022cmc: This is a jetted TDE (Andreoni et al. 2022; Pasham et al. 2023), similar to Sw J1644 + 57, which unlike the other jetted TDEs, was discovered in optical wavelengths.
AT 2022agi: This is a repeating TDE (see Subsection 4.3), also known as F01004. The first flare was reported in Tadhunter et al. (2017), while the second was studied in Sun et al. (2024).
AT 2020ksf: This TDE was first reported in Gilfanov et al. (2020), where a soft X-ray transient source was found to be coincident with AT 2020ksf. Furthermore, Alexander et al. (2021) reported faint radio emission detections coincident with the objects position.
AT 2017gge: This TDE was first reported in Wang et al. (2022) (ATLAS17jrp) and later appeared in Onori et al. (2022). Recently, it was reported to be the second TDE associated with a high-energy neutrino in Li et al. (2024), following AT 2019dsg (Stein et al. 2021b).
AT 2017eqx: This TDE was first reported in Nicholl et al. (2019), with broad H i and H ii emission.
LSQ12dyw: This TDE debuted first in a few circulars (Inserra et al. 2012; Smartt et al. 2012; Reis et al. 2012), where its nature was discussed. It was later studied and included in the sample of Charalampopoulos et al. (2022).
PTF09djl, PTF09ge: These TDEs were first presented in Arcavi et al. (2014), where they were characterized as TDE candidates. Later, they were included in the TDE sample of Charalampopoulos et al. (2022).
PS1-10jh: This TDE first appeared in Gezari et al. (2012), where its spectroscopic signature showed broad He ii emission lines. It was later included in the Charalampopoulos et al. (2022) TDE sample.
iPTF15af: This TDE first appeared in Blagorodnova et al. (2019), with broad He ii emission in its optical spectrum and several other broad features in the UV spectrum. It was later included in the Charalampopoulos et al. (2022) TDE sample.
iPTF16axa: This TDE was introduced as a candidate in Hung et al. (2017), where it showed broad hydrogen and helium emission lines. It was later included in the Charalampopoulos et al. (2022) TDE sample.
ASASSN-14li: This TDE was discovered in December of 2014 (Jose et al. 2014) and has since been studied extensively across the electromagnetic spectrum, in X-rays (Miller et al. 2015; Holoien et al. 2016b; Pasham et al. 2017), in the optical and near-ultraviolet (UV) (Cenko et al. 2016; Holoien et al. 2016b; Pasham et al. 2017), infrared (IR) (Jiang et al. 2016) and radio (Alexander et al. 2016; van Velzen et al. 2016). It was later included in the Charalampopoulos et al. (2022) TDE sample.
ASASSN-14ae: This TDE was discovered in January of 2014 (Prieto et al. 2014) and was first studied in Holoien et al. (2014) as a candidate, where the initial spectrum showed broad hydrogen emission and later evolved into having both broad helium and hydrogen emission. This source was also included in the Charalampopoulos et al. (2022) TDE sample.
ASASSN-15oi: This TDE was discovered in August of 2015 (Brimacombe et al. 2015) and was studied in Holoien et al. (2016a), where its optical spectrum showed broad helium features. It was later included in the Charalampopoulos et al. (2022) TDE sample.
Swift J1644+57: First, this source was thought to be a long lasting -ray outburst, discovered by the Swift BAT instrument. It was later revealed to be a TDE (Burrows et al. 2011; Bloom et al. 2011), since it decayed almost following the TDE characteristic power law. Moreover, Zauderer et al. (2011) reported radio detections coincident with this source, shortly after its discovery. The properties of this source led to the conclusion that Swift J1644+57 is a highly beamed, non-thermal, relativistic, jetted X-ray TDE.
Swift J2058+08: This source was detected shortly after the discovery of Swift J1644+57, once again by Swift BAT, sharing many similarities with it. It was first reported in Krimm et al. (2011) and further studied in Cenko et al. (2012), where a multi-wavelength follow-up was initiated. Cenko et al. (2012) detected a radio counterpart to the flare, suggesting that Swift J2058+08 is the second non-thermal (relativistic) jetted X-ray TDE.
Appendix B X-ray data reduction
B.1 Swift-XRT
B.1.1 Data Reduction
The XRT photon counting (PC) mode data were downloaded from HEASARC161616https://heasarc.gsfc.nasa.gov/ data archive, and processed using the XRTDAS software (Capalbi et al. 2005) developed at the ASI Science Data Center and included in the HEAsoft package (v. 6.33.2) distributed by HEASARC, using a procedure similar to that illustrated in Paggi et al. (2013).
For each observation calibrated and cleaned PC mode event files were produced with the xrtpipeline task (ver. 0.13.7), also producing exposure maps for each observation. In addition to the screening criteria used by the standard pipeline processing, we applied a further filter to screen background spikes that can occur when the angle between the pointing direction of the satellite and the bright Earth limb is low. In order to eliminate this so called bright Earth effect, due to the scattered optical light that usually occurs towards the beginning or the end of each orbit, we used the procedure proposed by Puccetti et al. (2011) and D’Elia et al. (2013). We monitored the count-rate on the CCD border and, through the xselect package, we excluded time intervals when the count-rate in this region exceeded . In addition we selected only time intervals with CCD temperatures less than (instead of the standard limit of ) since contamination by dark current and hot pixels, which increase the low energy background, is strongly temperature dependent (D’Elia et al. 2013).
B.1.2 Source Detection and Flux estimates
To detect X-ray sources in the XRT images, we made use of the ximage detection algorithm detect, which locates the point sources using a sliding-cell method. The average background intensity is estimated in several small square boxes uniformly located within the image. The position and intensity of each detected source are calculated in a box whose size maximizes the signal-to-noise ratio.
For each XRT-PC observation, we considered as coordinates of the X-ray counterpart to the TDE source the coordinates of the detected XRT-PC source closest to the TDE source coordinates if this happens to lie closer than . If no XRT-PC source was detected closer than to the TDE source coordinates, we used the TDE source coordinates themselves.
We then evaluated net count-rates (or their upper limits) at X-ray counterpart coordinates with the sosta algorithm that, besides the net count-rates and the respective uncertainties, yields the statistical significance of each source. In addition, sosta also estimates the optimal extraction radius that maximizes the signal to noise ratio of the source. We note that we used count-rates produced by sosta rather than those given by detect because the former are in most cases more accurate, since detect uses a global background for the entire image, whereas sosta uses a local background.
In order to get a first estimate of fluxes for TDE sources X-ray counterparts, we extracted appropriate arf and rmf files at each source location, making use of the xrtproducts task. As extraction regions we used circles centred at the X-ray source coordinates and with a radius equal to . Assuming black body model with a temperature of and an absorption component fixed to the Galactic value (Kalberla et al. 2005), we then converted the net count-rates evaluated earlier in intrinsic (i.e. unabsorbed) fluxes.
B.1.3 Spectral Extraction
To obtain better estimates on the X-ray source fluxes - as well as possible spectra variability - we extracted XRT-PC source spectra for the selected X-ray counterparts to TDE sources.
In general, source spectra - with the corresponding arf and rmf files - were extracted form events with xrtproducts task, using circular regions centred at the X-ray source coordinates with radii equal , while background spectra were extracted from annuli centred at the X-ray source coordinates, with inner and outer radii equal to and , respectively, excluding nearby detected sources.
When the source count-rate is above , the data are significantly affected by pileup in the inner part of the PSF (Moretti et al. 2005). To remove the pile-up contamination, we extract only events contained in an annular region centred on the X-ray source coordinates (Perri et al. 2007). While the outer radius of the annulus was set at , the inner radius was determined by comparing the observed profiles with the updated XRT PSF analytical model171717https://www.swift.ac.uk/analysis/xrt/SWIFT-XRT-CALDB-10_v01.pdf.
B.2 Chandra-ACIS
B.2.1 Data Reduction
Chandra-ACIS data were retrieved from the Chandra Data Archive181818http://cda.harvard.edu/chaser, and were processed and analysed with the Chandra Interactive Analysis of Observations (CIAO, Fruscione et al. 2006) data analysis system version 4.16 and Chandra calibration database CALDB version 4.11.1, adopting standard procedures. We run the ACIS level 2 processing with chandra_repro to apply up-to-date calibrations (CTI correction, ACIS gain, bad pixels), and then excluded time intervals of background flares exceeding with the deflare task. We produced full-band exposure maps, psf maps, and pileup maps with the pileup_map task.
B.2.2 Source Detection and Flux estimates
We run the wavdetect task to identify point sources in each observation with a sequence of wavelet scales (i.e., 1 1.41 2 2.83 4 5.66 8 11.31 16 pixels) and a false-positive probability threshold of .
As we did for Swift-XRT (see Sect. B.1.2), we considered as coordinates of the X-ray counterpart to the TDE source the coordinates of the detected Chandra-ACIS source closest to the TED source coordinates if this happens to lie closer than . If no Chandra-ACIS source was detected closer than to the TDE source coordinates, we extracted counts at the TDE source coordinates.
To evaluate the source fluxes we made use of the srcflux task, extracting counts from circular regions centered at the source location, with radii equal to the 99% EEF PSF radius (as estimated from the PSF maps), while for the background regions we considered annuli with inner and outer radii equal to and , respectively,excluding nearby detected sources. Appropriate arf and rmf files were generated for the considered regions, and the net count-rates were converted to intrinsic fluxes assuming a model comprising an absorption component fixed to the Galactic value and a black body with a temperature of . In the case of unconstrained count-rates, upper limits were evaluated with the aprates_tool task, and converted to intrinsic flux upper limits with the modelflux task, adopting the same model described before.
B.2.3 Spectral Extraction
Source spectra and the corresponding arf and rmf files were extracted form event files with specextract task, using circular regions centred at the X-ray source coordinates with radii equal , while background spectra were extracted from annuli centred at the X-ray source coordinates, with inner and outer radii equal to and , respectively, excluding nearby detected sources. Also in this case we excluded inner pixels with pileup larger than as estimated from the pileup maps.
B.3 XMM-Newton-EPIC
B.3.1 Data Reduction
XMM-Newton-EPIC data were retrieved from the XMM-Newton Science Archive191919http://nxsa.esac.esa.int/nxsa-web and reduced with the SAS202020http://www.cosmos.esa.int/web/xmm-newton/sas 21.0.0 software.
Following Nevalainen et al. (2005) we filtered EPIC data for hard-band flares by excluding the time intervals where the (for MOS1 and MOS2) or (for PN) count-rate evaluated on the whole detector FOV was more than 3 away from its average value. To achieve a tighter filtering of background flares, we iteratively repeated this process two more times, re-evaluating the average hard-band count-rate and excluding time intervals away more than from this value. The same procedure was applied to soft band restricting the analysis to an annulus with inner and outer radii of and excluding sources in the field, where the detected emission is expected to be dominated by the background.
B.3.2 Source Detection and Flux estimates
When possible, we merged data from MOS1, MOS2 and PN detectors using the emosaic task, in order to detect the fainter sources that wouldn’t be detected otherwise. Sources were detected following the standard SAS sliding box task edetect_chain that mainly consist of three steps: 1) source detection with local background, with a minimum detection likelihood of 8; 2) remove sources in step 1 and create a smooth background maps by fitting a 2-D spline to the residual image; 3) source detection with the background map produced in step 2 with a minimum detection likelihood of 10. The task emldetect was then used to determine the parameters for each input source - including the count-rate - by means of a maximum likelihood fit to the input images, selecting sources with a minimum detection likelihood of 15 and a flux in the band larger than (assuming an energy conversion factor of ). An analytical model of the PSF was evaluated at the source position and normalized to the source brightness. The source extent was then evaluated as the radius at which the PSF level equals half of local background.
Again, we considered as coordinates of the X-ray counterpart to the TDE source the coordinates of the detected XMM-Newton-EPIC source closest to the TDE source coordinates if this happens to lie closer than . If no XMM-Newton-EPIC source was detected closer than to the TDE source coordinates, we considered as coordinates of the possible X-ray counterpart the TDE source coordinates themselves. In this case, the source net count-rate (or its upper limits) and optimal extraction radius (maximizing the signal to noise ratio of the source) was provided by eregionanalyse.
Again, to estimate the fluxes of TDE sources X-ray counterparts, we extracted appropriate arf and rmf files for each available detector at each source location, making use of the arfgen and rmfgen tasks. As extraction regions we used circles centred at the X-ray source coordinates and with a radius equal to . Assuming a black body model with a temperature of and a absorption component fixed to the Galactic value, we then converted the net count-rates of each detector intrinsic (i.e. unabsorbed) fluxes. The mean source flux was then finally evaluated as the mean of the fluxes of each available detector weighted by its uncertainty, and the uncertainty on this mean flux was evaluated as:
where the sum runs on the available detectors (Webb et al. 2020).
B.3.3 Spectral Extraction
The source spectra were extracted with the evselect task from circular regions centred at the X-ray source coordinates with radii equal , while background spectra were extracted from annuli centred at the X-ray source coordinates, with inner and outer radii equal to and , respectively, excluding nearby detected sources. The corresponding arf and rmf files were generated with the rmfgen and arfgen tasks to take into account time and position-dependent EPIC responses. Again, inner regions of high pileup were estimated and excluded using the epatplot task through the distortion of pattern distribution, following the procedure explained in the SAS Data Analysis Threads212121https://www.cosmos.esa.int/web/xmm-newton/sas-thread-epatplot.
B.4 X-Ray Spectral Fitting
Spectral fitting was performed with the Sherpa222222http://cxc.harvard.edu/sherpa modelling and fitting application (Freeman et al. 2001) in the energy range for Swift-XRT and XMM-Newton-EPIC spectra, and in the energy range for Chandra-ACIS spectra, adopting Gehrels weighting (Gehrels 1986). Source spectra were binned to a minimum of 20 counts/bin to ensure the validity of statistics. In addition, for the EPIC spectra we excluded from the spectral fitting the band due to variable Al K lines, and fitted simultaneously MOS1, MOS2 and PN spectra, if available.
For the spectral fitting we used two different models:
1.
a model comprising an absorption component fixed to the Galactic value and a power-law with slope
2.
a model comprising an absorption component fixed to the Galactic value and a black body with temperature
In order to select the best fit model we made use of the Akaike information criterion (AIC, see for example Liddle 2007):
where is the number of model parameters, so we selected as best fit model the one that provides the lower value of AIC.
B.5 X-ray Catalogs
B.5.1 XMM Catalogue
We cross-matched the 4XMM-DR13 catalog with our catalogue of TDEs using a search radius. Since the fluxes reported in 4XMM-DR13 are derived from the count-rates assuming a model comprising a power-law with slope 1.7 and a Galactic absorption of (Webb et al. 2020), for uniformity we converted the count-rates for each detector to intrinsic fluxes using PIMMS232323https://heasarc.gsfc.nasa.gov/docs/software/tools/pimms.html, assuming a model comprising an absorption component fixed to the Galactic value and a black body with a temperature of . The fluxes for each detector and their respective uncertainties were there converted in mean flux and uncertainty using the same procedure illustrated in Sect. B.3.2.
B.5.2 EROSITA
We cross-matched the eRASS1 catalogue with our catalogue of TDEs using a search radius The fluxes reported in eRASS1 are derived from the count-rates assuming a model comprising a power-law with slope 2.0 and a Galactic absorption of (Merloni et al. 2024), therefore we converted the count-rates reported in the catalogue to intrinsic fluxes using PIMMS, again assuming a model comprising an absorption component fixed to the Galactic value and a black body with a temperature of .
For the TDE sources for which we did not find an eRASS1 counterpart, we obtained eRASS1 upper limits242424https://erosita.mpe.mpg.de/dr1/AllSkySurveyData_dr1/UpperLimitServer_dr1/. The upper limit eRASS fluxes were again converted in fluxes using PIMMS.
Appendix C Archival repeating flares
C.1 AT 2022dbl
This source was proposed to be a pTDE by Lin et al. (2024). The first flare was observed in 2022, while the second in 2024, with a 1.9 year interval. The light curve for this TDE is plotted in the upper panel of Fig. 6(c). Lin et al. (2024) argue that due to the similar optical spectra of the two flares, more specifically the broad Balmer, N iii and possible He ii emission lines, the most probably scenario is the repeating pTDE one. We also compare the shape of the two flares in the bottom panel of Fig. 6(c) and find that they are similar.
Moreover, Lin et al. (2024) tried to structure the orbit of the star around the SMBH and found, under the assumption of a BH of mass equal to , an elliptical orbit with period 710 days (time interval between the two flares) and a solar-like star, the eccentricity to be 0.997. Nonetheless, the pTDE nature of this phenomenon will be confirmed if a third flare is observed, since the orbital period of this event is calculated to be relatively short (1.9 years).
C.2 AT 2022agi
As briefly mentioned in Sect. 2, this TDE was studied by Sun et al. (2024), showcasing two flares. The first flare was observed in 2010, while the second in 2021, with a time interval of years (Sun et al. 2024). The light curve for this source is plotted in the upper panel of Fig. 6(d). Similarly to AT 2024pvu, the first flare was observed by the CRTS survey. The second outburst was observed by both the ATLAS and Gaia surveys, in multiple wavebands. Sun et al. (2024) discuss in great detail the nature of the repeating flares; whether it is a pTDE, double TDE (binary star system disrupted by the SMBH) or two separate events. In the end they are not able to exclude any of the aforementioned categories, with all being theoretically possible.
C.3 AT 2021mhg
This source was first discovered to be a TDE spectroscopically in 2021, while it was later reported to have a re-brightening in 2023 (Munoz-Arancibia et al. 2023). However, Somalwar et al. (2023b) attribute the second flare to a supernova (SN) Type Ia, that happened at the same position as the original 2021 TDE (see their Appendix A). Somalwar et al. (2023b) found that the 2023 light curve was fast-evolving and red, along with the fact that the optical spectra obtained 36 days after the peak was consistent with Type Ia SN observed roughly 30 days post-peak. To the best of our knowledge, this is the first and only event where a TDE was followed by a SN in the same position. As discussed by Somalwar et al. (2023b), this scenario should be taken into account when identifying pTDEs and repeating TDEs.
The light curves for the two flares are plotted in Fig. 6(f). It is evident that the shape of the outbursts is very different. This is more evident in the bottom panel of Fig. 6(f), where the two flares are completely inconsistent with one another. The peak of the SN outburst is much sharper than the TDE one. Moreover, the rise and decay time scales are much faster in the former compared to the latter.
C.4 AT 2020vdq
This TDE was part of the Yao et al. (2023) sample. The first flare was detected in 2020, while the second flare was observed in 2023 (Charalampopoulos et al. 2023). Somalwar et al. (2023b) studied AT 2020vdq’s repeating flare, which also appeared in the sample of Somalwar et al. (2023a), where a radio counterpart detection was reported, 1.4 years after the 2020 flare peak. There is no available spectrum from the time of the first flare, however the spectrum taken during the 2023 flare is consistent with the TDE-He class. Somalwar et al. (2023b) conclude that only the pTDE scenario is likely, excluding the scenario that the two flares could be caused by two separate events.
However, they also state that the host galaxy of AT 2020vdq is a post-starburst galaxy (see their Table 8), meaning that the rate of TDEs occurring are boosted. Using the TDE rate in Yao et al. (2023), which is , and the fact that in post-starburst galaxies the TDE rate is enhanced by one or two factors of magnitude, Somalwar et al. (2023b) give a probability of 30% of two independent TDEs occurring in the same galaxy from the 33 TDEs in the Yao et al. (2023) sample. This leads to the conclusion that the two flares observed in AT 2020vdq could have been caused by two separate events.
Additionally, the two flares have different rise times and decay times. This disparity is clear in the bottom panel of Fig. 7(b). The second flare has a sharper rise and decay time, drawing similarities to the case of AT 2021mhg. Moreover, the second peak is more than two times higher than that of the first flare. There is also cooling during the second flare, while the first doesn’t show much evolution in the temperature (Somalwar et al. 2023b).
C.5 AT 2019azh
This TDE shows a lot of similarities with AT 2024pvu, in the sense that the initial flare was also observed by the CRTS survey, while the second was detected by another optical survey, in this case ZTF and Gaia surveys, roughly 14.5 years later. The peak for the first flare was not recorded, probably due to the seasonal gap. Nonetheless, the peak of the second flare was observed in March of 2019. Hinkle et al. (2021) have reported the existence of the first flare, but as they discuss, there is only photometric data available from one wave band, which cannot confirm or deny whether the outburst was caused by a TDE or some other transient. This is the same as for AT 2024pvu, where the initial flare was observed only in one waveband and no spectrum was taken. Hinkle et al. (2021) also placed the two flares on top of each other (see their Figure 3). Although there is limited data for the first flare, the two flares seem to decay in a similar manner, further showcasing the resemblance between repeating flares.
However, unlike AT 2024pvu, there are a lot of information about the host galaxy of AT 2019azh, KUG 0810+227. Hinkle et al. (2021) were able to rule out strong AGN activity using archival data of KUG 0810+227. Nevertheless, they do not exclude the presence of a weak or low-luminosity AGN (LLAGN; Tozzi et al. 2006; Marchesi et al. 2016; Liu et al. 2017; Ricci et al. 2017). Moreover, they classify the host as a post-starburst galaxy, which as we previously discussed for AT 2020vdq, has boosted TDE rates, meaning that the initial flare could have indeed been a separate TDE.
C.6 AT 2018fyk
This TDE was first studied in Wevers et al. (2019). It remained X-ray and UV bright for at least a period of 500 days post discovery. There was another detection in the X-rays, 1216 days after the initial discovery (Wevers et al. 2023), although not as bright as the first flare (factor lower than the first flare). Wevers et al. (2023) attributed the re-brightening to the partial disruption of a star that was a part of a binary star system. There is also a notable bump in the optical light curve obtained by the Gaia survey (see Fig. 11(a)). The first outburst was observed in 2018, while the re-brightening was detected in 2021. Similar behaviour can be observed in Fig. 11(b), where we plot the catalogue X-ray data (see Sect. B.5).
Appendix D Repeating flare timescale comparison
To quantitatively investigate the similarity of the flares in the repeating flare sample, we estimate the rise times, decay times, peak flux and duration (metrics) for the non-confirmed TDE flares as well (i.e. first flare of AT2019teq, first flare of AT2022agi, first flare of AT2019azh, first flare of AT2024pvu, second flare of AT2021mhg). We note that we were unable to compute the rise time of AT2021mhg’s second flare and the decay time of the first flare of AT2019teq due to substantial gaps in the respective light curves. Additionally, we exclude AT2018fyk because the second flare is very faint in the optical band. For AT2022exr and AT2021uvz, which are the double peak flares, as well as AT2020acka, we cannot evaluate the shape of their flares since they are overlapping.
Figure 12 shows the metrics of the first flare vs the metrics of the second flare. Almost in all cases (except AT2019teq rise time and AT2024pvu in all timescales), the repeating flares are consistent within uncertainty. This was expected for AT2019teq, since the rise of the first flare is much faster than the second flare (TDE).